# Online Boosting Adaptive Learning under Concept Drift for Multistream Classification

### 17 Dec 2023 v1

@@ Introduction
To address all issues in the multi-stream classification task, we propose the Online Boosting Adaptive Learning (OBAL) method. As shown in Figure S1, OBAL consists of two stages, the first of which is the initialization phase, where we propose the AdaCOSA algorithm. The fundamental principle of AdaCOSA involves an adaptive interaction between models learned in the original source space and those acquired in the target space, aiming to align the temporal covariate shift and explore the dynamic relationships between different data streams based on feedback from the target domain. This process reinforces positive knowledge transfer, leading to optimal model migration. The second stage involves the online processing phase, during which our primary aim is to detect and adapt to the asynchronous drift in each data stream in real time. To achieve this, we employ the Drift Detection Method (DDM) (Gama et al. 2004) for labeled source streams, as it offers a stable and accurate detection approach. Simultaneously, we utilize the Gaussian Mixture Model (GMM) (Oliveira, Minku, and Oliveira 2021) based weighting strategy for asynchronous drift adaptation in these streams. For the unlabeled target stream, we design two sliding windows and continuously monitor their distribution changes to effectively detect drift occurrences. Once a drift is detected in the target stream, it signifies that the dynamic relationships learned in the first stage are no longer applicable, necessitating a return to the first stage for reinitialization. The main contributions of our work can be summarized as follows:
@@
Multi-source-stream Classification. It involves $N$ labeled source streams $S=\left\{S_{1},S_{2},\cdots,S_{N}\right\}$ and one unlabeled target stream $T$. Each arrived data sample at time $t$ is represented by $S_{i}(\boldsymbol{x}_{t},y_{t})$, where $\boldsymbol{x}_{t}\in\mathcal{D}^{d}$ is the $d$ features, and $y_{t}$ is the true label of the instance which can only be observed in $S_{i}$, $i\in\{1,2,\cdots,N\}$. It aims to build a classification model to predict the class label of $T(\boldsymbol{x}_{t})$ using the $S_{i}(\boldsymbol{x}_{t},y_{t})$ and $T(\boldsymbol{x}_{t})$.
@@
0:  Labels predicted on ${T}$ data.
@@
3:  Create the source detectors and adaptors $DDM_{Si}$, $GMM_{Si}$
4:  Create the target Gaussian Mixture Model $GMM_{T}$.
5:  Create the detection and reference windows ${W}_{det}$, ${W}_{ref}$.
@@
15:     Move detention window and calculate $\mu_{\text{det}},\mu_{\text{ref }}$.
@@
17:        Remove all base classifiers and return to line 2.
@@
As stated in Challenge 3, asynchronous concept drifts may occur in either the source or target streams over time. Therefore, for any given stream, it is necessary to continuously monitor its drifting situation in real time and promptly perform drift adaptation to accommodate the new concept.
Source stream processing:
For scenarios involving source drift, existing supervised drift detectors such as DDM can be employed, which offers more accurate drift detection because of the leveraging of labels. As a new source sample ${S}_{i}(\boldsymbol{x}_{t})$ arrives, the source classifier predicts its label, and then the drift detector is updated based on the prediction error. If no drift is detected, we will incrementally train the target classifier using the weighted mapped ${S}^{*}_{i}(\boldsymbol{x^{*}}_{t},y)$ with its corresponding weight $cw_{si}^{t}$. Since we have obtained the optimal weights $\mathbf{cw}_{Si}=[cw_{Si}^{1},cw_{Si}^{2},\cdots cw_{Si}^{L_{n}}]$ for the archived data batch during the initialization stage, we can retrieve the most relevant data from the archived data batch and assign its weights to the new coming data by indexing the minimum $L_{2}$ distance between new coming and archived data instances.
However, once a drift is detected within each source stream, an adaptation module should be deployed to handle new concepts. Here, we utilize the GMM to evaluate the distributions of the old and new concepts. GMM assumes several mixture components can model all real-world data, and it is formulated as follows:
@@
where $K$ represents the total number of Gaussians or mixture components, and $x$ is the observed multivariate. $w_{k}$ is a weight that is determined by the observations that constitute $C_{k}$, and $10\leq w_{k}\leq 1,\sum_{k=1}^{K}w_{k}=$ 1. $P\left(x\mid C_{k}\right)$ represents the likelihood of observation $x$ being assigned to mixture component $C_{k}$. It can be calculated by using the mean and the covariance of each mixture component $C_{k}$ as follows:
@@
Target stream processing:
In order to detect the drift in the target stream without utilizing labels, we use the archived target data batch $D_{T}$ to initialize a GMM model and deploy two sliding windows to detect the changes over time. Specifically, we design two sliding windows, i.e., Reference Window $W_{ref}=\left\{T\left(x_{1}\right),\cdots,T\left(x_{n}\right)\}\right.$ and Detect Window $W_{det}=\left\{T\left(x_{n+1}\right),\cdots,T\left(x_{2n}\right)\}\right.$, where $n$ is the instance number within the window and it is set as $n=L_{n}$. Then, the average conditional probability of the reference window can be calculated by a point estimation of the mean for the normal distribution,
@@
Otherwise, $W_{ref}$ and $W_{det}$ move step by step to receive new incoming data, i.e., $W_{ref}=\left\{T\left(x_{i}\right),\cdots,T\left(x_{n+i-1}\right)\right\}$ and $W_{det}=\left\{T\left(x_{n+i}\right),\cdots,T\left(x_{2n+i-1}\right)\right\}$. Once a change is detected, the historical base classifier becomes ineffective for classifying the target samples. Consequently, all base classifiers are eliminated from the base classifier pool, and the model undergoes re-initialization to adapt to the new concepts. The whole learning process is summarized in Algorithm 2, and a more detailed description is shown in Supplementary S2.
@@
In the experiment, we first empirically demonstrated that OBAL consistently outperforms current methods in multi-stream classification, highlighting both robustness and superiority. Second, we validated the substantial impact of dynamic inter-stream relationships on prediction, emphasizing the effectiveness of the AdaCOSA by ablation study. Finally, we established the scalability of OBAL across diverse data streams, corroborating its stable predictive capabilities. In addition, we also analyzed parameter sensitivity and the time complexity and cost of the algorithm.
Datasets Drift types Type #Instances #Features #Classes
Synthetic SEA Sudden/recurring Single 25K * 4 3 2
Tree Sudden/gradual Single 5K * 4 20 2
RBF Incremental Single 5K * 4 10 2
Hyperplane Incremental Single 30K* 4 4 2
Real-world Weather Unknown Single 4.5K* 4 8 2
Kitti Unknown Single 6.25K * 4 55 8
CNNIBN Unknown Multistream 30K * 4 124 2
BBC Unknown Multistream 30K * 4 124 2
Table 1: Characteristics of all datasets including 3 sources and 1 target stream.
@@
SEA 85.04±${}_{\pm 0.84}$ 85.78±${}_{\pm 0.92}$ 84.31±${}_{\pm 1.13}$ 88.42±${}_{\pm 1.70}$ 88.74±${}_{\pm 1.75}$ 87.62±${}_{\pm 1.01}$ 89.18±${}_{\pm 0.77}$ 90.23±${}_{\pm 1.42}$ 87.46±${}_{\pm 2.12}$ 90.98±${}_{\pm 0.87}$
Tree 76.98±${}_{\pm 1.11}$ 76.74±${}_{\pm 1.00}$ 75.21±${}_{\pm 1.07}$ 76.43±${}_{\pm 2.17}$ 76.71±${}_{\pm 1.86}$ 76.07±${}_{\pm 2.42}$ 78.93±${}_{\pm 0.61}$ 76.87±${}_{\pm 3.47}$ 77.64±${}_{\pm 1.47}$ 78.45±${}_{\pm 1.01}$
RBF 82.03±${}_{\pm 1.41}$ 83.46±${}_{\pm 1.20}$ 81.03±${}_{\pm 1.73}$ 84.53±${}_{\pm 2.01}$ 85.21±${}_{\pm 1.85}$ 83.16±${}_{\pm 2.13}$ 86.04±${}_{\pm 0.39}$ 85.26±${}_{\pm 2.89}$ 86.26±${}_{\pm 0.77}$ 86.78±${}_{\pm 0.91}$
Hyperplane 83.29±${}_{\pm 0.67}$ 84.05±${}_{\pm 0.52}$ 82.17±${}_{\pm 0.57}$ 86.17±${}_{\pm 1.04}$ 87.07±${}_{\pm 1.21}$ 86.01±${}_{\pm 1.49}$ 86.38±${}_{\pm 0.57}$ 87.66±${}_{\pm 1.74}$ 84.04±${}_{\pm 1.42}$ 88.01±${}_{\pm 1.17}$
Weather 71.04±${}_{\pm 1.50}$ 70.65±${}_{\pm 1.32}$ 72.17±${}_{\pm 1.17}$ 74.57±${}_{\pm 1.94}$ 75.03±${}_{\pm 2.01}$ 74.62±${}_{\pm 1.77}$ 77.74±${}_{\pm 0.89}$ 76.55±${}_{\pm 1.41}$ 76.02±${}_{\pm 3.43}$ 79.22±${}_{\pm 2.07}$
Kitti 54.21±${}_{\pm 2.61}$ 52.36±${}_{\pm 2.72}$ 50.38±${}_{\pm 2.43}$ 52.78±${}_{\pm 3.78}$ 54.01±${}_{\pm 3.09}$ 53.26±${}_{\pm 3.21}$ 50.29±${}_{\pm 1.34}$ 67.79±${}_{\pm 3.16}$ 64.82±${}_{\pm 4.17}$ 70.29±${}_{\pm 3.42}$
CNNIBN 66.76±${}_{\pm 0.74}$ 67.54±${}_{\pm 1.11}$ 65.34±${}_{\pm 0.92}$ 62.78±${}_{\pm 1.44}$ 65.74±${}_{\pm 1.76}$ 62.65±${}_{\pm 1.38}$ 68.79±${}_{\pm 0.31}$ 69.07±${}_{\pm 1.40}$ 68.83±${}_{\pm 0.89}$ ${}_{\pm 0.77}$
BBC 61.76±${}_{\pm 0.09}$ 61.26±${}_{\pm 0.43}$ 59.86±${}_{\pm 0.19}$ 62.78±${}_{\pm 1.16}$ 62.34±${}_{\pm 0.83}$ 60.76±${}_{\pm 0.77}$ 68.04±${}_{\pm 0.01}$ 63.36±${}_{\pm 1.07}$ 60.12±${}_{\pm 1.51}$ 66.43±${}_{\pm 1.42}$
Benchmark Datasets
As shown in Table 1, we conduct the experiment on four synthetic datasets (i.e., SEA (Street and Kim 2001), Tree (Liu, Lu, and Zhang 2020), RBF (Song et al. 2021a), and Hyperplane (Bifet and Gavalda 2007) ) and four popular real-world datasets (Weather (Ditzler and Polikar 2012), Kitti (Geiger, Lenz, and Urtasun 2012), CNNIBN (Vyas et al. 2014), and BBC (Vyas et al. 2014)), and more detailed descriptions of each dataset and multistream scenario simulation can be found in Supplementary S3.
Baselines
@@
This experiment first investigates whether using multi-source streams improves predictive capability compared to a single-source stream. From Figure 2, we can observe that the performance of multi-source streams outperforms single-stream performance on all datasets. This indicates that multi-source streams can provide additional complementary information to enhance predictive performance.
However, as the number of source streams increases, there may be a decline in performance. For example, the performance with five source streams is better than that with seven source streams on the Tree dataset. This may be because as the number of source streams increases, the complexity of the model also increases, which affects its performance. Overall, the performance of OBAL is stable across various sources, which demonstrates that our proposed method can easily adapt to different numbers of data streams.
@@
Time complexity and execution time
@@ Appendix A Supplementary
31:        remove all base classifiers and return to line 2.
@@
SEA 50.07 52.45 49.83 52.31 51.47 50.07 3.01 51.30 43.79 16.67
Tree 37.76 36.13 35.87 40.42 39.65 39.43 4.17 37.79 41.04 14.93
RBF 32.57 33.14 33.76 43.01 42.35 42.50 3.04 43.10 35.04 16.87
Hyperplane 57.41 55.73 56.20 52.38 53.09 53.42 4.10 54.75 57.17 19.37
Weather 60.41 59.14 57.23 52.01 51.49 52.37 4.97 51.70 89.43 10.71
Kitti 71.23 71.75 70.93 94.87 96.21 95.43 10.43 48.42 77.85 45.11
Table S1: Execution time (s) of Various Methods on All Benchmarks.
@@
$L_{n}$ $I_{max}$ $|P|$
SEA 200 3 5
Tree 200 3 5
RBF 300 4 10
Hyperplane 400 3 5
Weather 100 4 5
Kitti 100 5 2
CNNIBN 200 3 5
BBC 300 3 10
Table S2: Parameter settings on different datasets.

### 01 Jan 2024 v2

@@ Introduction
To address all issues in the multi-stream classification task, we propose the Online Boosting Adaptive Learning (OBAL) method. As shown in Figure S1, OBAL consists of two stages, the first of which is the initialization phase, where we propose the AdaCOSA algorithm. The fundamental principle of AdaCOSA involves an adaptive interaction between models learned in the original source space and those acquired in the target space, aiming to align the temporal covariate shift and explore the dynamic relationships between different data streams based on feedback from the target domain. This process reinforces positive knowledge transfer, leading to optimal model migration. The second stage involves the online processing phase, during which our primary aim is to detect and adapt to the asynchronous drift in each data stream in real-time. To achieve this, we employ the Drift Detection Method (DDM) (Gama et al. 2004) for labeled source streams, as it offers a stable and accurate detection approach. Simultaneously, we utilize the Gaussian Mixture Model (GMM) (Oliveira, Minku, and Oliveira 2021) based weighting strategy for asynchronous drift adaptation in these streams. For the unlabeled target stream, we design two sliding windows and continuously monitor their distribution changes to effectively detect drift occurrences. Once a drift is detected in the target stream, it signifies that the dynamic relationships learned in the first stage are no longer applicable, necessitating a return to the first stage for reinitialization. The main contributions of our work can be summarized as follows:
@@
Multi-source-stream Classification. It involves $N$ labeled source streams $S=\left\{S_{1},S_{2},\cdots,S_{N}\right\}$ and one unlabeled target stream $T$. Each arrived data sample at time $t$ is represented by $S_{i}(\boldsymbol{x}_{t},y_{t})$, where $\boldsymbol{x}_{t}\in\mathcal{D}^{d}$ is the $d$-dimensional features, and $y_{t}$ is the true label of the instance which can only be observed in $S_{i}$, $i\in\{1,2,\cdots,N\}$. It aims to build a classification model to predict the class label of $T(\boldsymbol{x}_{t})$ using the $S_{i}(\boldsymbol{x}_{t},y_{t})$ and $T(\boldsymbol{x}_{t})$.
@@
0:  Source streams $\left\{{S}_{1},{S}_{2},\cdots,{S}_{N}\right\}$, target stream ${T}$, classifier pool $P$, initial sample size $L_{n}$.
0:  Predicted labels for target stream.
@@
3:  Create $DDM_{Si}$ and $GMM_{Si}$ for source stream.
4:  Create $GMM_{T}$ for target stream.
5:  Create detection and reference windows ${W}_{det}$, ${W}_{ref}$.
@@
15:     Move detection window and calculate $\mu_{\text{det}},\mu_{\text{ref }}$.
@@
17:        Remove all base classifiers and return to line 1.
@@
SEA Tree RBF Hyperplane Weather Kitti CNNIBN BBC
FUSIONs1 85.04±${}_{\pm 0.84}$ 76.98±${}_{\pm 1.11}$ 82.03±${}_{\pm 1.41}$ 83.29±${}_{\pm 0.67}$ 71.04±${}_{\pm 1.50}$ 54.21±${}_{\pm 2.61}$ 66.76±${}_{\pm 0.74}$ 61.76±${}_{\pm 0.09}$
FUSIONs2 85.78±${}_{\pm 0.92}$ 76.74±${}_{\pm 1.00}$ 83.46±${}_{\pm 1.20}$ 84.05±${}_{\pm 0.52}$ 70.65±${}_{\pm 1.32}$ 52.36±${}_{\pm 2.72}$ 67.54±${}_{\pm 1.11}$ 61.26±${}_{\pm 0.43}$
FUSIONs3 84.31±${}_{\pm 1.13}$ 75.21±${}_{\pm 1.07}$ 81.03±${}_{\pm 1.73}$ 82.17±${}_{\pm 0.57}$ 72.17±${}_{\pm 1.17}$ 50.38±${}_{\pm 2.43}$ 65.34±${}_{\pm 0.92}$ 59.86±${}_{\pm 0.19}$
ATLs1 88.42±${}_{\pm 1.70}$ 76.43±${}_{\pm 2.17}$ 84.53±${}_{\pm 2.01}$ 86.17±${}_{\pm 1.04}$ 74.57±${}_{\pm 1.94}$ 52.78±${}_{\pm 3.78}$ 62.78±${}_{\pm 1.44}$ 62.78±${}_{\pm 1.16}$
ATLs2 88.74±${}_{\pm 1.75}$ 76.71±${}_{\pm 1.86}$ 85.21±${}_{\pm 1.85}$ 87.07±${}_{\pm 1.21}$ 75.03±${}_{\pm 2.01}$ 54.01±${}_{\pm 3.09}$ 65.74±${}_{\pm 1.76}$ 62.34±${}_{\pm 0.83}$
ATLs3 87.62±${}_{\pm 1.01}$ 76.07±${}_{\pm 2.42}$ 83.16±${}_{\pm 2.13}$ 86.01±${}_{\pm 1.49}$ 74.62±${}_{\pm 1.77}$ 53.26±${}_{\pm 3.21}$ 62.65±${}_{\pm 1.38}$ 60.76±${}_{\pm 0.77}$
Melanie 89.18±${}_{\pm 0.77}$ 78.93±${}_{\pm 0.61}$ 86.04±${}_{\pm 0.39}$ 86.38±${}_{\pm 0.57}$ 77.74±${}_{\pm 0.89}$ 50.29±${}_{\pm 1.34}$ 68.79±${}_{\pm 0.31}$ 68.04±${}_{\pm 0.01}$
AOMSDA 90.23±${}_{\pm 1.42}$ 76.87±${}_{\pm 3.47}$ 85.26±${}_{\pm 2.89}$ 87.66±${}_{\pm 1.74}$ 76.55±${}_{\pm 1.41}$ 67.79±${}_{\pm 3.16}$ 69.07±${}_{\pm 1.40}$ 63.36±${}_{\pm 1.07}$
MCMO 87.46±${}_{\pm 2.12}$ 77.64±${}_{\pm 1.47}$ 86.26±${}_{\pm 0.77}$ 84.04±${}_{\pm 1.42}$ 76.02±${}_{\pm 3.43}$ 64.82±${}_{\pm 4.17}$ 68.83±${}_{\pm 0.89}$ 60.12±${}_{\pm 1.51}$
OBAL (ours) 90.98±${}_{\pm 0.87}$ 78.45±${}_{\pm 1.01}$ 86.78±${}_{\pm 0.91}$ 88.01±${}_{\pm 1.17}$ 79.22±${}_{\pm 2.07}$ 70.29±${}_{\pm 3.42}$ ${}_{\pm 0.77}$ 66.43±${}_{\pm 1.42}$
@@
As stated in Challenge 3, asynchronous concept drifts may occur in either the source or target streams over time. Therefore, for any given stream, it is necessary to continuously monitor its drifting situation in real-time and promptly perform drift adaptation to accommodate the new concept.
Source Stream Processing.
For scenarios involving source drift, existing supervised drift detectors such as DDM can be employed, which offers more accurate drift detection because of the leveraging of labels. As a new source sample ${S}_{i}(\boldsymbol{x}_{t})$ arrives, the source classifier predicts its label, and then the drift detector is updated based on the prediction error. If no drift is detected, we will incrementally train the target classifier using the weighted mapped ${S}^{*}_{i}(\boldsymbol{x^{*}}_{t},y)$ with its corresponding weight $cw_{si}^{t}$. Since we have obtained the optimal weights $\mathbf{cw}_{Si}=[cw_{Si}^{1},cw_{Si}^{2},\cdots cw_{Si}^{L_{n}}]$ for the archived data batch during the initialization stage, we can retrieve the most relevant data from the archived data batch and assign its weights to the new coming data by indexing the minimum $L_{2}$ distance between new coming and archived data instances. However, once a drift is detected within each source stream, an adaptation module should be deployed to handle new concepts. Here, we utilize the GMM to evaluate the distributions of the old and new concepts. GMM assumes several mixture components can model all real-world data, and it is formulated as follows:
@@
where $K$ represents the total number of Gaussians or mixture components, and $x$ is the observed multivariate. $w_{k}$ is a weight that is determined by the observations that constitute $C_{k}$, and $10\leq w_{k}\leq 1,\sum_{k=1}^{K}w_{k}=$ 1. $P\left(x\mid C_{k}\right)$ represents the likelihood of observation $x$ being assigned to mixture component $C_{k}$. It can be calculated by using the mean $\mu_{k}$ and the covariance $\Sigma_{k}$ of each mixture component $C_{k}$:
@@
Target Stream Processing.
To detect the drift in the target stream without utilizing labels, we use the archived target data batch $D_{T}$ to initialize a GMM model and deploy two sliding windows to detect the changes over time. Specifically, we design two sliding windows, i.e., Reference Window $W_{ref}=\left\{T\left(x_{1}\right),\cdots,T\left(x_{n}\right)\}\right.$ and Detect Window $W_{det}=\left\{T\left(x_{n+1}\right),\cdots,T\left(x_{2n}\right)\}\right.$, where $n$ is the instance number within the window and it is set as $n=L_{n}$. Then, the average conditional probability of the reference window can be calculated by a point estimation of the mean for the normal distribution,
@@
Otherwise, $W_{ref}$ and $W_{det}$ move step by step to receive new incoming data, i.e., $W_{ref}=\left\{T\left(x_{t}\right),\cdots,T\left(x_{n+t-1}\right)\right\}$ and $W_{det}=\left\{T\left(x_{n+t}\right),\cdots,T\left(x_{2n+t-1}\right)\right\}$. Once a change is detected, the historical base classifier becomes ineffective for classifying target samples. Consequently, all base classifiers are eliminated from the classifier pool, and the model undergoes re-initialization to adapt to the new concepts. The learning process is summarized in Algorithm 2.
@@
In the experiment, we first empirically demonstrated that OBAL consistently outperforms current methods, highlighting both robustness and superiority. Second, we validated the substantial impact of dynamic inter-stream relationships on prediction, emphasizing the effectiveness of the AdaCOSA by ablation study. Additionally, we confirmed OBAL’s scalability across various data streams, validating its consistent predictive performance. Finally, we assessed parameter sensitivity, time complexity, and algorithmic cost.
@@
Benchmark Datasets.
We conduct the experiment on four synthetic datasets (i.e., SEA (Street and Kim 2001), Tree (Liu, Lu, and Zhang 2020), RBF (Song et al. 2021a), and Hyperplane (Bifet and Gavalda 2007) ) and four popular real-world datasets (Weather (Ditzler and Polikar 2012), Kitti (Geiger, Lenz, and Urtasun 2012), CNNIBN (Vyas et al. 2014), and BBC (Vyas et al. 2014)), and more detailed descriptions of each dataset and multistream scenario simulation can be found in Supplementary S3 and Table S1.
Baselines.
@@
This experiment first investigates whether using multi-source streams improves predictive capability compared to a single-source stream. From Figure 2, we can observe that the performance of multi-source streams outperforms single-stream performance on all datasets. This indicates that multi-source streams can provide additional complementary information to enhance predictive performance. However, as the number of source streams increases, there may be a decline in performance. For example, the performance with five source streams is better than that with seven source streams on the Tree dataset. This may be because as the number of source streams increases, the complexity of the model also increases, which affects its performance. Overall, the performance of OBAL is stable across various sources, which demonstrates that our proposed method can easily adapt to different numbers of data streams.
@@
Time Complexity and Execution Time.
@@ Appendix A Supplementary
Datasets Drift types Type #Instances #Features #Classes
Synthetic SEA Sudden/recurring Single 25K * 4 3 2
Tree Sudden/gradual Single 5K * 4 20 2
RBF Incremental Single 5K * 4 10 2
Hyperplane Incremental Single 30K* 4 4 2
Real-world Weather Unknown Single 4.5K* 4 8 2
Kitti Unknown Single 6.25K * 4 55 8
CNNIBN Unknown Multistream 30K * 4 124 2
BBC Unknown Multistream 30K * 4 124 2
Table S1: Characteristics of all datasets including 3 sources and 1 target stream.
@@
31:        remove all base classifiers and return to line 1.
@@
$L_{n}$ $I_{max}$ $|P|$
SEA 200 3 5
Tree 200 3 5
RBF 300 4 10
Hyperplane 400 3 5
Weather 100 4 5
Kitti 100 5 2
CNNIBN 200 3 5
BBC 300 3 10
Table S2: Parameter settings on different datasets.
@@
SEA 50.07 52.45 49.83 52.31 51.47 50.07 3.01 51.30 43.79 16.67
Tree 37.76 36.13 35.87 40.42 39.65 39.43 4.17 37.79 41.04 14.93
RBF 32.57 33.14 33.76 43.01 42.35 42.50 3.04 43.10 35.04 16.87
Hyperplane 57.41 55.73 56.20 52.38 53.09 53.42 4.10 54.75 57.17 19.37
Weather 60.41 59.14 57.23 52.01 51.49 52.37 4.97 51.70 89.43 10.71
Kitti 71.23 71.75 70.93 94.87 96.21 95.43 10.43 48.42 77.85 45.11
Table S3: Execution time (s) of Various Methods on All Benchmarks.